Three charges lie along the x-axis. the positive charge q1=15µC is at x=2.0m, and the positive charge q2=6.0µC is at origin. Where must a negative charge q3 must be placed on the x axis so that the resultant electric force on it is zero?
1
Expert's answer
2019-05-17T11:52:39-0400

According to Coulomb’s law the force F between two point charges, q1 and q2, separated by a distance r is equal to


"F=k\frac{\left| {{q}_{1}}{{q}_{2}} \right|}{{{r}^{2}}}"

q1=15µC, q2=6.0µC, r=2m. Let r13 and r23 are the distances between q1 and q3 as well as between q2 and q3, respectively, r13 + r23=r=2.



For net force to be zero on q3 must be performed

"{{\vec{F}}_{net}}={{\vec{F}}_{13}}+{{\vec{F}}_{23}}=0"

or

"{{\vec{F}}_{13}}=-{{\vec{F}}_{23}}"

For absolute values of forces we get

"k\frac{{{q}_{1}}{{q}_{3}}}{r_{13}^{2}}=k\frac{{{q}_{2}}{{q}_{3}}}{r_{23}^{2}}"

or

"\frac{{{q}_{1}}}{{{\left( r-{{r}_{23}} \right)}^{2}}}=\frac{{{q}_{2}}}{r_{23}^{2}}"

Solve this equation for   r23

"\frac{{{\left( r-{{r}_{23}} \right)}^{2}}}{r_{23}^{2}}=\frac{{{q}_{1}}}{{{q}_{2}}}\,\,\,\Rightarrow \,\,\,\frac{r-{{r}_{23}}}{{{r}_{23}}}=\sqrt{\frac{{{q}_{1}}}{{{q}_{2}}}}"

Then

"{{r}_{23}}=\frac{r}{1+\sqrt{\frac{{{q}_{1}}}{{{q}_{2}}}}}"

Substitute known values

"{{r}_{23}}=\frac{2\,m}{1+\sqrt{\frac{15\mu C}{6.0\mu C}}}=0.77\,meters"

Since the charge q2 is at the origin, then the coordinate of charge q3 is x3=0.77meters